OpenStudy (anonymous):

Find the x-values (if any) at which f is not contininous. Which of the discontinuities are removable?

OpenStudy (anonymous):

Are you adding a post with the function f?

OpenStudy (anonymous):

\[f(x)= 1/3x + 1 x \le2\]

OpenStudy (anonymous):

gah, typing tihs completely is confusing. sorry. one moment please

OpenStudy (anonymous):

No problem, go ahead.

OpenStudy (anonymous):

f(x) = 1/3x + 1 when x is less than or equal to 2 and f(x)= 3-x when x is greater than 2

OpenStudy (anonymous):

i understand there is a discontinuity at 2, but how can I tell if its removable or non-removable?

OpenStudy (anonymous):

Oh, good... you're better off than I thought :)

OpenStudy (anonymous):

Except I guess I don't have a good definition of "removable" discontinuity

OpenStudy (anonymous):

The fact that the values of f(x) for x less than 2 then "jump" (the discontinuity) when x changes to be equal or greater than 2 makes me think you can't "remove" that somehow, but it might depend on the actual definition of "removable"

OpenStudy (anonymous):

:/ still confused